Lucy's Cookies sells 4,800 biscuit cases per year. Biscuits are produced at a rate of 1,000 cases of biscuits per day. The case cost is $ 26.5 . Holding cost is 20% of the case cost. Set-up cost is $ 11,000 . Calculate the Economic Production Quantity (EPQ) in cases. Your Answer:
Answer #1

EPQ = V(2 X DXS) H(1-d/p) = V(2 X 4800 X 2.29) 5.3(1-13/1000) = V(21984) 5.23 = V(21984) 5.2311 = 65 cases per day Explanatio